a notebook has an area of 150m² and a perimeter of 50cm. what are the dimensions of the notebook
Found 2 solutions by Theo, greenestamps:
Answer by Theo(13342)

l * w = 150
2l + 2w = 50
solve for l in the first equation to get:
l = 150/w
replace l in the second equation with that to get:
2*150/w + 2w = 50
simplify to get:
300/w + 2w = 50
multiply both sides of the equation by w to get:
300 + 2w^2 = 50w
divide both sides of the equation by 2 to get:
150 + w^2 = 25w
subtract 25w from both sides of the equation and reorder the terms in descenng order of degree to get:
w^2 - 25w + 150 = 0
factor to get:
(w-10) * (w-15) = 0
solve for w to get:
w = 10 or w = 15
when w = 10, l = 150/10 = 15
when w = 15, l = 150/15 = 10
dimensions of the notebook are 10 by 15 or 15 by 10, however you want to look at it.


Answer by greenestamps(13200)



The perimeter is 50, so length plus width is 25.

x = width
25-x = length

The area is 150:

x(25-x)=150
25x-x^2=150
x^2-25x=-150
x^2-25x+150=0
(x-15)(x-10)=0

ANSWER: The dimensions of the notebook are 10cm and 15cm.